# Joint Value Estimation and Bidding in Repeated First-Price Auctions

### v1

@@ 1 Introduction
However, another important challenge in FPAs often overlooked by the existing literature is the estimation of the actual value of an impression (or advertising slot). In practice, this value is determined by the difference between two potential outcomes—commonly clicks or conversion rates—when the auction is won versus lost. As shown in Fig. 1, an advertiser can observe the outcome of winning the advertising slot when he/she wins the auction, and the outcome of doing nothing when losing the auction; the value of the advertising slot is the difference between these two potential outcomes. Consequently, as [34] observes, the problem of value estimation can be framed as a specific instance of causal inference, where the target quantity is the treatment effect, and the treatment is determined by the bidding outcome. This introduces the following challenges:
@@ 1.2 Organization
The rest of the paper is organized as follows. Section 2 introduces the problem formulation and presents the main regret bounds under three different models. Sections 3, 4 and 5 work on the three models of adversarial outcomes, linear potential outcomes, and linear treatment effect, respectively, and develop the corresponding bidding strategies under each scenario. Conclusions and some future directions are discussed in Section 6. The proofs of technical results, including the regret lower bounds, are deferred to the appendices.
@@ 1.3 Notation
For positive integers $n\in\mathbb{N}$, we let $[n]=\{1,2,\dots,n\}$. For an event $E$, we defined the indicator function $\mathds{1}[E]$ to be 1 if the event occurs and 0 otherwise. The standard asymptotic notations $O(\cdot),\Omega(\cdot),$ and $\Theta(\cdot)$ are used to suppress constant factors, and ${\widetilde{O}}(\cdot),\widetilde{\Omega}(\cdot),$ and $\widetilde{\Theta}(\cdot)$ to suppress poly-logarithmic factors. For a random variable $X$, ${\mathbb{E}}[X]$ and $\mathsf{Var}(X)$ denote its mean and variance respectively. Throughout, we use the filtration ${\mathcal{F}}_{t}$ to denote all available history up to the beginning of round $t$, and write ${\mathbb{E}}_{t}[X]={\mathbb{E}}[X|{\mathcal{F}}_{t}]$ for simplicity. For a vector $x\in{\mathbb{R}}^{d}$ and a PSD matrix $A\in{\mathbb{R}}^{d\times d}$, let $\|x\|_{A}=\sqrt{x^{\top}Ax}$. For a cumulative distribution function $G$ on $[0,1]$, let its inverse be $G^{-1}(z)=\inf\{x\in[0,1]:G(x)\geq z\}$.
@@ 2 Problem Formulation and Main Results
In this section, we formulate this learning problem in FPAs and the three models under different assumptions on the potential outcomes. Subsequently, we discuss the key challenges of these formulations in Section 2.2, and summarize our main results in Section 2.3.
@@ 2.1 Problem Formulation
As in many causal inference problems, Assumption 2.1 of no unmeasured confounders is vital to avoid Simpson’s paradox and make estimators like inverse propensity weighting (IPW) work. Specializing to bidding in FPAs, even for $v_{t,0}=0$ and $x_{t}=\varnothing$, dependence between $v_{t,1}$ and $M_{t}$ could lead to a $\Omega(T)$ regret (see, e.g. [24, Theorem 8] and [15, Theorem 8]). As for the joint distribution of $(v_{t,1},v_{t,0},x_{t})$ over time, we also adopt the following conditional independence assumption, which essentially states that the context sequence $\{x_{t}\}_{t=1}^{T}$ is chosen by an oblivious adversary (e.g. $x_{t}$ cannot depend on $b_{t-1}$ or $M_{t-1}$).
@@
Thanks to Assumption 2.1, Assumption 2.2, and the full-information feedback on $M_{t}$, the estimation of $M_{t}$ is done independently of either value estimation or bidding: regardless of the generation processes of $(v_{t,1},v_{t,0},b_{t})$, the bidder always observes the same trajectory $\{(x_{t},M_{t})\}_{t=1}^{T}$, which is also the only useful information for predicting future HOBs. By contrast, the problem of joint value estimation and bidding lies in the interdependence between $(v_{t,1},v_{t,0})$ and $b_{t}$: the bidder chooses the bid $b_{t}$ based on the current estimate of $v_{t,1}-v_{t,0}$, while the choice of $b_{t}$ affects the bidder’s likelihood of observing the potential outcomes $v_{t,1}$ or $v_{t,0}$. Since the primary focus of this paper is on the problem of joint value estimation and bidding, in the sequel we will focus less on the modeling and estimation of $M_{t}$, and use the following black-box model instead.
@@ 2.1.1 Model 1: Adversarial Outcomes
In the first model, we have no context (i.e. $x_{t}=\varnothing$), and make no distributional assumptions on the potential outcomes $(v_{t,1},v_{t,0})$. Specifically, we assume that both $\{(v_{t,1},v_{t,0})\}_{t=1}^{T}$ and the CDFs $\{G_{t}\}_{t=1}^{T}$ (satisfying Assumption 2.3) are generated by an oblivious adversary. Thanks to the independence of $(v_{1,t},v_{0,t})$ and $M_{t}$ in Assumption 2.1, the expected payoff takes the form (2). In addition, since no context is available, a natural (context-dependent) oracle is to bid the optimal fixed price $b_{*}\in{\mathcal{B}}$ in hindsight, i.e. the regret of a policy $\pi=(b_{1},\dots,b_{T})$ in this model is defined as
@@ 2.2 Key Challenges
    Would the bidding process lead to a propensity score too close to 00 or $1$, and if so, incur a large estimation error and a large regret in future bidding?
@@
The optimal bid $b_{t}$, however, is often likely to either win or lose the auction—i.e. either $G_{t}(b_{t})$ or $1-G_{t}(b_{t})$ approaches 00. This leads to an unbounded variance in the estimator $\widehat{e}_{t}(b_{t})$ and thereby breaks the regret guarantee for linear contextual bandit algorithms in the literature. To address this issue, the common approach to reduce the estimation error is to impose the overlap condition, i.e. require that $G_{t}(b_{t})\in[\eta,1-\eta]$ for some $\eta>0$ and therefore lead to non-greedy strategies for $b_{t}$. Nonetheless, we show that the overlap condition is not necessary by exploiting two equivalent forms of the reward function in (2), with leading coefficient $G_{t}(b_{t})$ or $1-G_{t}(b_{t})$ on the decision $b_{t}$. Our final bidding algorithm then selects the better reward formulation in a data-driven manner, and manages to mitigate the effect of a possibly large variance in decision making.
@@ 2.3 Main Results
Under Assumptions 2.12.3 and Model 1 in Section 2.1.1, there is an algorithm $\pi_{\mathrm{adv}}$ that achieves the expected regret
@@
Based on Assumption 2.5, we have the following regret guarantee for Model 2.
@@
Under Assumptions 2.12.3, 2.5, and Model 2 in Section 2.1.2, there is an algorithm $\pi_{\mathrm{lpo}}$ that achieves the expected regret
@@
Theorem 2.6 shows an $\widetilde{O}(\sqrt{dT})$ regret whenever the estimation of the HOB distribution $G_{t}$ does not suffer from a too large error. For example, when $M_{t}$ is i.i.d., the famous Dvoretzky–Kiefer–Wolfowitz inequality [19] states that $\varepsilon_{t}=\widetilde{O}(1/\sqrt{t})$ for the empirical CDF $\widehat{G}_{t}$, and $\Delta=\widetilde{O}(\sqrt{T})$ is smaller than the $\widetilde{O}(\sqrt{dT})$ regret.
@@
Based on Assumption 2.7, we have the following regret guarantee for Model 3.
@@
Under Assumptions 2.12.3, 2.7, and Model 3 in Section 2.1.3, there is an algorithm $\pi_{\mathrm{lte}}$ (with the knowledge of the confidence bounds $\{\delta_{t}\}_{t=1}^{T}$) that achieves the expected regret
@@
Again, when $M_{t}$ is i.i.d., Bernstein’s concentration inequality shows that Assumption 2.7 holds with $\delta_{t}=\widetilde{O}(1/\sqrt{t})$, so that $\Delta_{1}=\widetilde{O}(\sqrt{T})$ and $\Delta_{2}=\widetilde{O}(1)$. Therefore, Theorem 2.8 still leads to an $\widetilde{O}(\sqrt{dT})$ regret even if only the treatment effect possesses a linear structure. We also remark that Assumption 2.3 crucially does not assume a density lower bound; therefore, it could happen that the overlap condition is violated (i.e. $G_{t}(b_{t})$ is close to 00 or $1$), but the regret upper bound in Theorem 2.8 remains unaffected.
@@
The lower bounds in Theorem 2.9 are established by classical information-theoretic arguments, and for completeness we include the proofs in the appendix. Table 1 presents a summary of the above results.
@@ 3 Model 1: Adversarial Outcomes
1 Input: time horizon $T$, finite bid set ${\mathcal{B}}^{\prime}\subseteq{\mathcal{B}}$, learning rate $\eta\in\left(0,\frac{1}{2}\right)$.
@@ 6      Make bid $b_{t}$, observe $M_{t}$, and build payoff estimator for every $b\in{\mathcal{B}}^{\prime}$:
To handle the continuous bid space ${\mathcal{B}}$, we consider a simple discretization =${\mathcal{B}}^{\prime}=\left\{\frac{k}{T}:k\in[T]\right\}$. Thanks to the Lipschitzness of the CDF $G_{t}$ in Assumption 2.3, we are able to approximate the optimal fixed bid in ${\mathcal{B}}$ with the optimal bid in the discretization, which leads to the following result.
Corollary 3.2 (Theorem 2.4).
Let =${\mathcal{B}}^{\prime}=\left\{\frac{k}{T}:k\in[T]\right\}\subseteq{\mathcal{B}}$. Let $\pi$ be the resulting bidding strategy of Algorithm 1 running on this discretization with $\eta=\sqrt{\frac{\log(T)}{(4e-2)T}}$. Then under Assumptions 2.12.3, the policy $\pi$ achieves expected regret
@@ 4.1.1 Base algorithm
1 Input: Time indices $\Phi_{t}\subseteq[t-1]$, bid space $B_{t}\subseteq{\mathcal{B}}$, estimated CDF $\widehat{G}_{t}$, time horizon $T$.
@@
9       Compute the reward estimator $\widehat{r}_{t}\leftarrow\widehat{G}_{t}(b)\left(\hat{\theta}_{t,1}^{\top}x_{t}-b\right)+\left(1-\widehat{G}_{t}(b)\right)\hat{\theta}_{t,0}^{\top}x_{t}$.
@@
The expressions of $(\widehat{\theta}_{t,0},\widehat{\theta}_{t,1})$ are computed in Line 7 of Algorithm 2. After that, for each active bid $b\in B_{t}$, the base algorithm uses the plug-in approach to estimate the expected payoff function ${\mathbb{E}}[r_{t}(b)]$ in (2) by $\widehat{r}_{t}(b)$ in Line 9, as well as a width parameter $w_{t}(b)$ in Line 10. The rational behind these quantities is summarized in the following lemma.
@@ 4.1.2 Master algorithm
1 Input: Time horizon $T$, bid space ${\mathcal{B}}=[0,1]$, estimation oracle ${\mathcal{O}}$.
2Initialize: Set $S\leftarrow\lceil\log_{2}\sqrt{T}\rceil$, $\Phi^{(s)}_{1}=\varnothing$ for $s\in[S]$.
@@
14                   Choose $b_{t}\leftarrow\max_{i\in B_{s}}\widehat{r}^{(s)}_{t}(i)$.
@@
Note that the conditional independence assumption in Lemma 4.1 breaks down when $\Phi_{t}=[t-1]$ for all $t\in[T]$. To see this, note that the selection of the bid $b_{t}$ depends on the realized outcomes
@@
therefore, conditioned on such a future bid $b_{t}$, the past potential outcomes $(v_{s,0},v_{s,1})_{s<t}$ may become dependent. To address this subtle dependence issue, we propose to use a master algorithm SUP-LIN-PO in Algorithm 3. The idea behind this master algorithm is taken from [5], and has been subsequently used to address similar dependence issues in [16, 24]. The purpose of this master algorithm is to partition the time indices $[t]$ into multiple stages, eliminate the suboptimal bids in a hierarchical manner, and select the bid $b_{t}$ only through the confidence width $w_{t}(b)$ in LIN-PO, instead of the point estimates $\widehat{r}_{t}(b)$. Crucially, the widths $w_{t}(b)$ are computed based on the contexts $(x_{t},b_{t},M_{t})$ on the current stage, but do not involve the potential outcomes $(v_{t,1},v_{t,0})$ on this stage. By contrast, the point estimates $\widehat{r}_{t}(b)$ (which depend on the potential outcomes $(v_{s,1},v_{s,0})_{s<t}$ on this stage) are only revealed when we decide to move to the next stage (Line 17). This information structure turns out to be the key idea to ensure the conditional independence in Lemma 4.1:
@@
Under Assumption 2.1 and Assumption 2.2, for every $s\in[S]$ and $t\in[T]$, $(v_{\tau,1},v_{\tau,0})_{\tau\in\Phi_{t}^{(s)}}$ are conditionally independent given $(x_{\tau},b_{\tau},M_{\tau})_{\tau\in\Phi_{t}^{(s)}}$.
Finally, we remark that while statistically Algorithm 3 leads to the target regret upper bound for the continuous bid space ${\mathcal{B}}=[0,1]$, for computational considerations one can discretize ${\mathcal{B}}$ to a fine enough resolution which does not affect the regret bound. In addition, Line 12 of Algorithm 3 ensures that we never reach $s=S+1$ in the inner for loop.

### v2

@@ 1 Introduction
However, another important challenge in FPAs often overlooked by the existing literature is the estimation of the actual value of an impression (or advertising slot). In practice, this value is determined by the difference between two potential outcomes—commonly clicks or conversion rates—when the auction is won versus lost. As shown in Fig.˜1, an advertiser can observe the outcome of winning the advertising slot when he/she wins the auction, and the outcome of doing nothing when losing the auction; the value of the advertising slot is the difference between these two potential outcomes. Consequently, as [34] observes, the problem of value estimation can be framed as a specific instance of causal inference, where the target quantity is the treatment effect, and the treatment is determined by the bidding outcome. This introduces the following challenges:
@@ 1.2 Organization
The rest of the paper is organized as follows. Section˜2 introduces the problem formulation and presents the main regret bounds under three different models. Sections˜3, 4 and 5 work on the three models of adversarial outcomes, linear potential outcomes, and linear treatment effect, respectively, and develop the corresponding bidding strategies under each scenario. Conclusions and some future directions are discussed in Section˜6. The proofs of technical results, including the regret lower bounds, are deferred to the appendices.
@@ 1.3 Notation
[…]
@@ 2 Problem Formulation and Main Results
In this section, we formulate this learning problem in FPAs and the three models under different assumptions on the potential outcomes. Subsequently, we discuss the key challenges of these formulations in Section˜2.2, and summarize our main results in Section˜2.3.
@@ 2.1 Problem Formulation
As in many causal inference problems, Assumption˜2.1 of no unmeasured confounders is vital to avoid Simpson’s paradox and make estimators like inverse propensity weighting (IPW) work. Specializing to bidding in FPAs, even for $v_{t,0}=0$ and $x_{t}=\varnothing$, dependence between $v_{t,1}$ and $M_{t}$ could lead to a $\Omega(T)$ regret (see, e.g. [24, Theorem 8] and [15, Theorem 8]). As for the joint distribution of $(v_{t,1},v_{t,0},x_{t})$ over time, we also adopt the following conditional independence assumption, which essentially states that the context sequence $\{x_{t}\}_{t=1}^{T}$ is chosen by an oblivious adversary (e.g. $x_{t}$ cannot depend on $b_{t-1}$ or $M_{t-1}$).
@@
Thanks to Assumption˜2.1, Assumption˜2.2, and the full-information feedback on $M_{t}$, the estimation of $M_{t}$ is done independently of either value estimation or bidding: regardless of the generation processes of $(v_{t,1},v_{t,0},b_{t})$, the bidder always observes the same trajectory $\{(x_{t},M_{t})\}_{t=1}^{T}$, which is also the only useful information for predicting future HOBs. By contrast, the problem of joint value estimation and bidding lies in the interdependence between $(v_{t,1},v_{t,0})$ and $b_{t}$: the bidder chooses the bid $b_{t}$ based on the current estimate of $v_{t,1}-v_{t,0}$, while the choice of $b_{t}$ affects the bidder’s likelihood of observing the potential outcomes $v_{t,1}$ or $v_{t,0}$. Since the primary focus of this paper is on the problem of joint value estimation and bidding, in the sequel we will focus less on the modeling and estimation of $M_{t}$, and use the following black-box model instead.
@@ 2.1.1 Model 1: Adversarial Outcomes
In the first model, we have no context (i.e. $x_{t}=\varnothing$), and make no distributional assumptions on the potential outcomes $(v_{t,1},v_{t,0})$. Specifically, we assume that both $\{(v_{t,1},v_{t,0})\}_{t=1}^{T}$ and the CDFs $\{G_{t}\}_{t=1}^{T}$ (satisfying Assumption˜2.3) are generated by an oblivious adversary. Thanks to the independence of $(v_{1,t},v_{0,t})$ and $M_{t}$ in Assumption˜2.1, the expected payoff takes the form (2). In addition, since no context is available, a natural (context-dependent) oracle is to bid the optimal fixed price $b_{*}\in{\mathcal{B}}$ in hindsight, i.e. the regret of a policy $\pi=(b_{1},\dots,b_{T})$ in this model is defined as
@@ 2.2 Key Challenges
    Would the bidding process lead to a propensity score too close to 0 or $1$, and if so, incur a large estimation error and a large regret in future bidding?
@@
The optimal bid $b_{t}$, however, is often likely to either win or lose the auction—i.e. either $G_{t}(b_{t})$ or $1-G_{t}(b_{t})$ approaches 0. This leads to an unbounded variance in the estimator $\widehat{e}_{t}(b_{t})$ and thereby breaks the regret guarantee for linear contextual bandit algorithms in the literature. To address this issue, the common approach to reduce the estimation error is to impose the overlap condition, i.e. require that $G_{t}(b_{t})\in[\eta,1-\eta]$ for some $\eta>0$ and therefore lead to non-greedy strategies for $b_{t}$. Nonetheless, we show that the overlap condition is not necessary by exploiting two equivalent forms of the reward function in (2), with leading coefficient $G_{t}(b_{t})$ or $1-G_{t}(b_{t})$ on the decision $b_{t}$. Our final bidding algorithm then selects the better reward formulation in a data-driven manner, and manages to mitigate the effect of a possibly large variance in decision making.
@@ 2.3 Main Results
Under Assumptions 2.12.3 and Model 1 in Section˜2.1.1, there is an algorithm $\pi_{\mathrm{adv}}$ that achieves the expected regret
@@
Based on Assumption˜2.5, we have the following regret guarantee for Model 2.
@@
Under Assumptions 2.12.3, 2.5, and Model 2 in Section˜2.1.2, there is an algorithm $\pi_{\mathrm{lpo}}$ that achieves the expected regret
@@
Theorem˜2.6 shows an $\widetilde{O}(\sqrt{dT})$ regret whenever the estimation of the HOB distribution $G_{t}$ does not suffer from a too large error. For example, when $M_{t}$ is i.i.d., the famous Dvoretzky–Kiefer–Wolfowitz inequality [19] states that $\varepsilon_{t}=\widetilde{O}(1/\sqrt{t})$ for the empirical CDF $\widehat{G}_{t}$, and $\Delta=\widetilde{O}(\sqrt{T})$ is smaller than the $\widetilde{O}(\sqrt{dT})$ regret.
@@
Based on Assumption˜2.7, we have the following regret guarantee for Model 3.
@@
Under Assumptions 2.12.3, 2.7, and Model 3 in Section˜2.1.3, there is an algorithm $\pi_{\mathrm{lte}}$ (with the knowledge of the confidence bounds $\{\delta_{t}\}_{t=1}^{T}$) that achieves the expected regret
@@
Again, when $M_{t}$ is i.i.d., Bernstein’s concentration inequality shows that Assumption˜2.7 holds with $\delta_{t}=\widetilde{O}(1/\sqrt{t})$, so that $\Delta_{1}=\widetilde{O}(\sqrt{T})$ and $\Delta_{2}=\widetilde{O}(1)$. Therefore, Theorem˜2.8 still leads to an $\widetilde{O}(\sqrt{dT})$ regret even if only the treatment effect possesses a linear structure. We also remark that Assumption˜2.3 crucially does not assume a density lower bound; therefore, it could happen that the overlap condition is violated (i.e. $G_{t}(b_{t})$ is close to 0 or $1$), but the regret upper bound in Theorem˜2.8 remains unaffected.
@@
The lower bounds in Theorem˜2.9 are established by classical information-theoretic arguments, and for completeness we include the proofs in the appendix. Table˜1 presents a summary of the above results.
@@ 3 Model 1: Adversarial Outcomes
1Input: time horizon $T$, finite bid set $\mathcal{B}$$^{\prime}\subseteq{\mathcal{B}}$, learning rate $\eta\in\left(0,\frac{1}{2}\right)$.
@@ 6  Make bid $b_{t}$, observe $M_{t}$, and build payoff estimator for every $b\in{\mathcal{B}}^{\prime}$:
To handle the continuous bid space $\mathcal{B}$}, we consider a simple discretization ${\mathcal{B}}^{\prime}=\left\{\frac{k}{T}:k\in[T]\right\}$. Thanks to the Lipschitzness of the CDF $G_{t}$ in Assumption 2.3, we are able to approximate the optimal fixed bid in $\mathcal{B}$} with the optimal bid in the discretization, which leads to the following result.
Corollary 3.2 (Theorem˜2.4).
Let ${\mathcal{B}}^{\prime}=\left\{\frac{k}{T}:k\in[T]\right\}\subseteq{\mathcal{B}}$. Let $\pi$ be the resulting bidding strategy of Algorithm˜1 running on this discretization with $\eta=\sqrt{\frac{\log(T)}{(4e-2)T}}$. Then under Assumptions 2.12.3, the policy $\pi$ achieves expected regret
@@ 4.1.1 Base algorithm
1Input: Time indices $\Phi_{t}\subseteq[t-1]$, bid space $B_{t}\subseteq{\mathcal{B}}$, estimated CDF $\widehat{G}_{t}$, time horizon $T$.
@@
9   Compute the reward estimator $\widehat{r}_{t}(b)\leftarrow\widehat{G}_{t}(b)\left(\hat{\theta}_{t,1}^{\top}x_{t}-b\right)+\left(1-\widehat{G}_{t}(b)\right)\hat{\theta}_{t,0}^{\top}x_{t}$.
@@
The expressions of $(\widehat{\theta}_{t,0},\widehat{\theta}_{t,1})$ are computed in Line 7 of Algorithm˜2. After that, for each active bid $b\in B_{t}$, the base algorithm uses the plug-in approach to estimate the expected payoff function ${\mathbb{E}}[r_{t}(b)]$ in (2) by $\widehat{r}_{t}(b)$ in Line 9, as well as a width parameter $w_{t}(b)$ in Line 10. The rational behind these quantities is summarized in the following lemma.
@@ 4.1.2 Master algorithm
1Input: Time horizon $T$, bid space ${\mathcal{B}}=[0,1]$, estimation oracle $\mathcal{O}$}.
2Initialize: Set $S\leftarrow\lceil\log_{2}\sqrt{T}\rceil$, $\Phi^{(s)}_{1}$$\varnothing$ for $s\in[S]$.
@@
14         Choose $b_{t}\leftarrow\operatorname*{arg\,max}_{i\in B_{s}}\widehat{r}^{(s)}_{t}(i)$.
@@
Note that the conditional independence assumption in Lemma˜4.1 breaks down when $\Phi_{t}=[t-1]$ for all $t\in[T]$. To see this, note that the selection of the bid $b_{t}$ depends on the realized outcomes
@@
therefore, conditioned on such a future bid $b_{t}$, the past potential outcomes $(v_{s,0},v_{s,1})_{s<t}$ may become dependent. To address this subtle dependence issue, we propose to use a master algorithm SUP-LIN-PO in Algorithm˜3. The idea behind this master algorithm is taken from [5], and has been subsequently used to address similar dependence issues in [16, 24]. The purpose of this master algorithm is to partition the time indices $[t]$ into multiple stages, eliminate the suboptimal bids in a hierarchical manner, and select the bid $b_{t}$ only through the confidence width $w_{t}(b)$ in LIN-PO, instead of the point estimates $\widehat{r}_{t}(b)$. Crucially, the widths $w_{t}(b)$ are computed based on the contexts $(x_{t},b_{t},M_{t})$ on the current stage, but do not involve the potential outcomes $(v_{t,1},v_{t,0})$ on this stage. By contrast, the point estimates $\widehat{r}_{t}(b)$ (which depend on the potential outcomes $(v_{s,1},v_{s,0})_{s<t}$ on this stage) are only revealed when we decide to move to the next stage (Line 17). This information structure turns out to be the key idea to ensure the conditional independence in Lemma˜4.1:
@@
Under Assumption˜2.1 and Assumption˜2.2, for every $s\in[S]$ and $t\in[T]$, $(v_{\tau,1},v_{\tau,0})_{\tau\in\Phi_{t}^{(s)}}$ are conditionally independent given $(x_{\tau},b_{\tau},M_{\tau})_{\tau\in\Phi_{t}^{(s)}}$.
Finally, we remark that while statistically Algorithm˜3 leads to the target regret upper bound for the continuous bid space ${\mathcal{B}}=[0,1]$, for computational considerations one can discretize $\mathcal{B}$} to a fine enough resolution which does not affect the regret bound. In addition, Line 12 of Algorithm˜3 ensures that we never reach $s=S+1$ in the inner for loop.